1250 person went to see a circus show each adult paid 75rupees and each child paid 25rupees for admission ticket find the number of adults and number of children if the total collection from them amounts to 61250rupees

Grade:10

1 Answers

Vikas TU
7 years ago
Let adults be x and children by y.
then,
x + y = 1250 …............(1)
and
75x + 25y = 61250.....................(2)
Solve the two eqns. and get x and y values.